# SingMOS-Pro: An Comprehensive Benchmark for Singing Quality Assessment

### v2

@@ Abstract
Singing voice generation progresses rapidly, yet evaluating singing quality remains a critical challenge. Human subjective assessment, typically in the form of listening tests, is costly and time consuming, while existing objective metrics capture only limited perceptual aspects. In this work, we introduce SingMOS-Pro, a dataset for automatic singing quality assessment. Building on our preview version SingMOS, which provides only overall ratings, SingMOS-Pro expands annotations of the additional part to include lyrics, melody, and overall quality, offering broader coverage and greater diversity. The dataset contains 7,981 singing clips generated by 41 models across 12 datasets, spanning from early systems to recent advances. Each clip receives at least five ratings from professional annotators, ensuring reliability and consistency. Furthermore, we explore how to effectively utilize MOS data annotated under different standards and benchmark several widely used evaluation methods from related tasks on SingMOS-Pro, establishing strong baselines and practical references for future research. The dataset can be accessed at https://huggingface.co/datasets/TangRain/SingMOS-Pro.
@@ 1 Introduction
In singing voice generation tasks, the conventional SQA method combines both human subjective assessment and objective assessments. Human subjective evaluation, typically in the form of Mean Opinion Score (MOS), is widely regarded as the ”gold standard” for SQA. However, conducting standard subjective tests is time-consuming and labor-intensive, and their results often lack comparability across different experiments. On the other hand, existing objective metrics, such as mel-cepstral distortion, have shown a weak correlation with perceived audio quality. This discrepancy makes the rapid and accurate evaluation of singing vocals challenging. Consequently, there is an urgent need for an efficient, reliable, and universal SQA method for singing voice generation.
@@ 2.1 Overview
For convenience, we define a system as the set of audio samples generated under a specific dataset–model–setting configuration. For convenience, we define a system as the set of audio samples generated under a specific dataset–model–setting configuration. In total, SingMOS-Pro contains 11.15 hours of audio, consisting of 7,981 mono singing clips and 44,247 ratings collected from 78 experienced annotators. The distribution of clips is illustrated in Fig. 1: the dataset includes 3,425 clips from 60 SVS systems (plus 250 clips from song generation systems), 1,307 clips from 17 SVC systems, 2,671 clips from 52 SVR systems, and 578 ground-truth recordings from 12 systems. Among them, 6,937 clips are resampled to 16kHz, while 631 and 413 remain at 24kHz and 44.1kHz, respectively. Across all samples, the average clip duration is 5.03 seconds. Each singing clip is evaluated by at least five experienced annotators for the overall MOS score. In addition, 4,155 clips are further annotated with lyrics scores and melody scores, which capture the clarity of pronunciation and the naturalness of melody, respectively.
@@ 2.2 Statistics
Fig. 2 shows the distribution of utterances in SingMOS-Pro with respect to overall MOS scores. Most generated singing vocals are concentrated around scores between 3 and 4, with fewer samples between 2 and 3. In contrast, ground-truth recordings are mainly clustered between 4 and 5, while only a few noisy samples appear near 1. Overall, the distribution approximates a Gaussian shape.
@@
Fig. 3: Distribution of Systems Across MOS Intervals.
@@ 3.1 Singing Clips Collection
Dataset: We collect 167 songs from the test sets of several publicly available singing datasets, including seven datasets for SVS/SVC from CtrSVDD [19], as well as additional three Mandarin (ACE-Opencpop [20], ACE-Kising [20], Chinese GTSinger [21]) and three Japanese datasets (e.g., Ameboshi Cipher Utagoe DB [^1] https://parapluie2c56m.wixsite.com/mysite, Natsume Singing [^2] https://github.com/AmanoKei/NatsumeSinging, Namine Ritsu Utagoe DB1). In addition, we construct an extra dataset by extracting melodies into musical scores with ChatMusician [22] and generating lyrics with DeepSeek-V2 [23].
Table 1: Models in SingMOS-Pro. Bold models indicate those evaluated under multiple settings. “X/Y” denotes the number of models X and the number of systems Y.
Category Models
@@
SVR (9/52) Codec: DAC [25], Encodec [26], SoundStream [27];
@@
SVS (26/60) OpenSource: RNN [1], XiaoiceSing [2], Diffsinger [10], VISinger [7], VISinger2 [8], VISinger2+ [9], TokSing [3], SingOMD [6], StyleSinger [11], TCSinger [32], TechSinger [33], Sinsy [18], ARNNSVS [18], DiffNNSVS [18], svsAug [34];
Demo: EveryoneCanSing [^3] https://everyone-can-sing.github.io/, ExpressiveSinger [^4] https://expressivesinger.github.io/ExpressiveSinger/, PeriodSinger [^5] https://rlataewoo.github.io/periodsinger/, SPSinger [^6] https://danny-nus.github.io/SPSinger/, TCSinger2777https://aaronz345.github.io/TCSinger2Demo/, XSinger [^8] https://jisang93.github.io/x-singer/;
Song Generation: ACEStep [35], Diffrhythm [36], Hailuo [^9] https://hailuoai.com/music, Suno [^10] https://suno.com/, Yue [37]
@@
In practice, part of the generated samples is directly obtained from the CtrSVDD Challenge 2024 [19], which follows the same procedure. For several models not publicly available (Demo in Table. 1), we supplement SingMOS-Pro with samples collected from their official demo pages. For song generation models, we synthesize samples with their demo models in the pop style using Chinese lyrics from Opencpop and GTSinger, and apply MelBand Roformer [42] to separate the vocals from the accompaniment.
In total, we collect 141 systems, producing 7,981 clips comprising both synthetic and real singing vocals. Fig. 4 illustrates the distribution of clips per system and the overall dataset duration.
@@ 3.2 Annotation Protocols
In SingMOS-Pro, we recruite 78 experienced annotators to conduct the evaluation. In total, 7,981 audio clips are annotated, yielding 44,247 overall-performance ratings. Among these, 4,155 clips are additionally annotated for lyric and melody, resulting in 23,475 lyric ratings and 23,475 melody ratings.
Evaluation Dimensions: The overall MOS focuses on the general performance of a singing clip. The lyrics MOS specifically evaluates the clarity and accuracy of pronunciation, while the melody MOS assesses the naturalness of the melody and whether the singing sounds harmonious.
Evaluation Design: All MOS tests are conducted online. After completing pre-annotation training, 78 annotators perform evaluations in quiet environments. The audio samples are distributed across five batches, with each clip annotated by at least five annotators. The first and forth batch collect only overall MOS, while the second, third and fifth batches include three dimensions: overall, lyrics, and melody scores. The fourth and fifth batches are annotated for other projects [34, 9]. All scores are collected on a 5-point Likert scale. Notably, the preview version SingMOS consists entirely of samples from the first batch.
@@ 3.3 Dataset Splits
Fig. 5: Distribution of Subset Across MOS Intervals.
To facilitate the effective use of our dataset, we predefine a train/test split. For the first, second, and third annotation batches, systems more than 50 clips are split into training and test sets with a 7:3 ratio. Systems containing 10–50 clips are entirely assigned to the test set, while systems with fewer than 10 clips are included only in the training set. For the fourth and fifth annotation batches, all results are merged into the training set. Since annotation standards vary across batches, we maintain separate test sets for different batches. The impact of these decisions on the training set will be further discussed in Section 4.2. After the split, the training set contains 4,453 clips, while test1, test2, and test3 consist of 1 070, 1 444, and 339 clips, respectively. The distribution of utterance-level MOS scores is illustrated in Fig. 5.
@@ 4.1 Experiments Setting
All experiments are conducted on SingMOS-Pro with only 16kHz audio, since our self-supervised learning (SSL) backbone supports only 16kHz inputs. All 24kHz and 44.1kHz samples are filtered out, resulting in 4,007 clips and 87 systems for training and 2,091/1,540/376 clips and 31/41/15 systems for test1/test2/test3. The backbone model is wav2vec 2.0 and is optimized with L1 loss with margin, using stochastic gradient descent with a learning rate of 0.001 and a momentum of 0.9. Training is performed for 200 epochs with a batch size of 15. For evaluation, we report Root Mean Squared Error (RMSE), Linear Correlation Coefficient (LCC), and Spearman’s Rank Correlation Coefficient (SRCC). Among these metrics, SRCC is considered the most important, since it reflects the ranking consistency of perceptual quality. To facilitate a comprehensive evaluation, we compute weighted averages of the results from test1, test2, and test3 at both the utterance and system levels, where the weights are determined by the corresponding number of utterances and systems in each test set.
@@ 4.2 Exploration on Train Set Utilization
In audio quality assessment, whether for speech or singing, it is challenging to collect a sufficiently large and diverse dataset under a unified annotation standard. This makes it crucial to effectively leverage data annotated with different criteria. Following the approaches in AlignNet [43] and SHEET [44], we investigate two strategies on a self-supervised learning backbone, where wav2vec 2.0 [45] is adopted. The first strategy is multi-dataset finetuning (MDF), in which we first train on the training set of the first batch for 10 epochs and then continue finetuning on the entire training set. The second strategy is the use of domain id.
As shown in Table 2, the plain SSL model without any auxiliary design and the SSL model equipped with both MDF and domain id achieve the best weighted SRCC at the system level and utterance level, respectively. Adding domain id leads to clear improvements on test3, suggesting that domain identifiers are particularly beneficial for test sets with fewer samples, as they help the model learn to mitigate domain discrepancies. Also, introducing MDF consistently improves performance on most metrics of test2, indicating that multi-dataset finetuning alleviates the confusion caused by different annotation standards across batches. Considering the weighted results across all test sets, the joint use of domain ids and MDF yields the most effective overall performance, while directly using the plain SSL model also proves to be a reasonable and convenient choice.
4.3 Exploration with Singing Quliaty Assessment
@@
Since SQA requires greater emphasis on melodic naturalness compared to speech assessment, we further explore several strategies for using pitch information. These strategies include the pitch histogram proposed in [46], as well as the direct use of MIDI pitch and MIDI pitch variance. As shown in the last three rows of Table 3, the pitch histogram yields slightly better performance than MIDI pitch, but the overall improvement over the SSL baseline remains marginal, underscoring the need for further exploration of how melodic cues can be more effectively integrated. Beyond these approaches, exploring how to incorporate melody scores and lyric scores into SQA constitutes a valuable avenue for future work.

### v3

@@ Abstract
Singing voice generation progresses rapidly, yet evaluating singing quality remains a critical challenge. Human subjective assessment, typically in the form of listening tests, is costly and time consuming, while existing objective metrics capture only limited perceptual aspects. In this work, we introduce SingMOS-Pro, a dataset for automatic singing quality assessment. Building on our preview version SingMOS, which provides only overall ratings, SingMOS-Pro expands annotations of the additional part to include lyrics, melody, and overall quality, offering broader coverage and greater diversity. The dataset contains 7,981 singing clips generated by 41 models across 12 datasets, spanning from early systems to recent advances. Each clip receives at least five ratings from experienced annotators, ensuring reliability and consistency. Furthermore, we explore how to effectively utilize MOS data annotated under different standards and benchmark several widely used evaluation methods from related tasks on SingMOS-Pro, establishing strong baselines and practical references for future research. The dataset can be accessed at https://huggingface.co/datasets/TangRain/SingMOS-Pro.
@@ 1 Introduction
In singing voice generation tasks, the conventional SQA method combines both human subjective assessment and objective assessments. Human subjective evaluation, typically in the form of Mean Opinion Score (MOS), is widely regarded as the ”gold standard” for SQA [21]. However, conducting standard subjective tests is time-consuming and labor-intensive, and their results often lack comparability across different experiments. On the other hand, existing objective metrics, such as mel-cepstral distortion, have shown a weak correlation with perceived audio quality [21, 19]. This discrepancy makes the rapid and accurate evaluation of singing vocals challenging. Consequently, there is an urgent need for an efficient, reliable, and universal SQA method for singing voice generation.
@@ 2.1 Overview
For convenience, we define a system as the set of audio samples generated under a specific dataset–model–setting configuration. In total, SingMOS-Pro contains 11.15 hours of audio, consisting of 7,981 mono singing clips and 44,247 ratings collected from 78 experienced annotators. The distribution of clips is illustrated in Fig. 1: the dataset includes 3,425 clips from 60 SVS systems (plus 250 clips from song generation systems), 1,307 clips from 17 SVC systems, 2,671 clips from 52 SVR systems, and 578 ground-truth recordings from 12 systems. Among them, 6,937 clips are resampled to 16kHz, while 631 and 413 remain at 24kHz and 44.1kHz, respectively. Across all samples, the average clip duration is 5.03 seconds. Each singing clip is evaluated by at least five experienced annotators for the overall MOS score. In addition, 4,155 clips are further annotated with lyrics scores and melody scores, which capture the clarity of pronunciation and the naturalness of melody, respectively.
@@ 2.2 Statistics
Fig. 2 shows the distribution of utterances in SingMOS-Pro with respect to overall MOS scores. Most generated singing vocals are concentrated around scores between 3 and 4, with fewer samples between 2 and 3. In contrast, ground-truth recordings are mainly clustered between 4 and 5, while only a few noisy samples appear near 1. Overall, the distribution approximates a bimodal shape.
@@
Fig. 3: Distribution of Systems Across MOS Intervals. Detailed of System ID is available at metadata in SingMOS-Pro.
@@ 3.1 Singing Clips Collection
Dataset: We collect 167 songs from the test sets of twelve publicly available singing datasets, covering over 50 singers. It includes five Mandarin datasets (Opencpop [31], M4Singer [42], ACE-Opencpop [25], Kising [25] and Chinese GTSinger [47]) and seven Japanese datasets (JVSMusic [27], Kiritan Singing [16], Namine Ritsu [^1] https://drive.google.com/drive/folders/1XA2cm3UyRpAk_BJb1LTytOWrhjsZKbSN, Ofuton P [^2] https://sites.google.com/view/oftn-utagoedb/%E3%83%9B%E3%83%BC%E3%83%A0, Ameboshi Cipher [^3] https://parapluie2c56m.wixsite.com/mysite, Natsume Singing [^4] https://github.com/AmanoKei/Natsume_Singing and Oniku Kurumi [^5] https://onikuru.info/db-download/ ). In addition, we construct a synthetic dataset by extracting melodies into musical scores with ChatMusician [38] and generating lyrics with DeepSeek-V2 [4].
Table 1: Models in SingMOS-Pro. Italicized models indicate those evaluated under multiple settings. “X/Y” denotes the number of models X and the number of systems Y.
Cate Models
@@
SVR(9/52) Codec: SoundStream [41], Encodec [1], DAC [20];
@@
SVS
(26/60)
Pretrained Model: VISinger2 [44], VISinger [43], XiaoiceSing [18], VISinger2+ [37], DiffNNSVS [36], RNN [22], Diffsinger [14], svsAug [48], StyleSinger [45], TCSinger [46], TechSinger [7], ACE Studio [25], SingOMD [29], TokSing [33], Sinsy [36], ARNNSVS [36];
Official Demo: ExpressiveSingera, SPSingerb, PeriodSingerc, EveryoneCanSingd, TCSinger2e, XSingerf;
Song Generation: ACEStep [6], Diffrhythm [15], Hailuog, Sunoh, Yue [39]
@@
In practice, part of the generated samples is directly obtained from the CtrSVDD Challenge 2024 [40], which follows the same procedure. For models not publicly available (Demo in Table. 1), we supplement SingMOS-Pro with samples collected from their official demo pages. For song generation models, we synthesize samples with their demo models in the pop style using Chinese lyrics from Opencpop and GTSinger, and apply MelBand Roformer [30] to separate the vocals from the accompaniment.
In total, we collect 141 systems, producing 7,981 clips comprising both synthetic and real singing vocals. Fig. 4 illustrates the distribution of clips per system and the duration of utterances.
@@ 3.2 Annotation Protocols
In SingMOS-Pro, we recruite 78 annotators to conduct the evaluation. In total, 7,981 audio clips are annotated, yielding 44,247 overall-performance ratings. Among these, 4,155 clips are additionally annotated for lyric and melody, resulting in 23,475 lyric ratings and 23,475 melody ratings.
Evaluation Dimensions: Lyrics MOS specifically measures pronunciation clarity and intelligibility, while Melody MOS evaluates melodic naturalness and pitch accuracy. Finally, Overall MOS provides a holistic assessment, capturing the general quality by considering both lyrical and melodic dimensions.
Evaluation Design: All MOS tests are conducted online. ensure data quality, all annotators are required to complete a pre-annotation training lecture. Evaluations are performed in quiet environments, and the audio samples are annotated in five discrete batches over time. Each audio clip receives at least five independent annotations. The first and forth batch collect only overall MOS, while the second, third and fifth batches include three dimensions: overall, lyrics, and melody scores. The fourth and fifth batches are annotated for other projects [48, 37]. All scores are collected on a 5-point Likert scale. Notably, the preview version SingMOS [28] consists entirely of samples from the first batch.
@@ 3.3 Dataset Splits
To facilitate the effective use of our dataset, we predefine a train/test split. For the first, second, and third annotation batches, systems more than 50 clips are split into training and test sets with a 7:3 ratio. Systems containing 10–50 clips are entirely assigned to the test set, while systems with fewer than 10 clips are included only in the training set. For the fourth and fifth annotation batches, all results are merged into the training set. Since annotation standards vary across batches, we maintain separate test sets for different batches. The impact of these decisions on the training set will be further discussed in Section 4.2. After the split, the training set contains 4,453 clips, while test1, test2, and test3 consist of 1 070, 1 444, and 339 clips, respectively.
@@ 4.1 Experiments Setting
All experiments are conducted on SingMOS-Pro with only 16kHz audio, since our self-supervised learning (SSL) backbone supports only 16kHz inputs. All 24kHz and 44.1kHz samples are filtered out, resulting in 4,007 clips and 87 systems for training and 2,091/1,540/376 clips and 31/41/15 systems for test1/test2/test3. The SSL backbone model is wav2vec2-large [2] and is optimized with L1 loss with margin, using stochastic gradient descent with a learning rate of 0.001 and a momentum of 0.9. Training is performed for 200 epochs with a batch size of 15. For evaluation, we report Root Mean Squared Error (RMSE), Linear Correlation Coefficient (LCC), and Spearman’s Rank Correlation Coefficient (SRCC). Among these metrics, SRCC is considered the most important, since it reflects the ranking consistency of perceptual quality. To facilitate a comprehensive evaluation, we compute weighted averages of the results from test1, test2, and test3 at both the utterance and system levels, where the weights are determined by the corresponding number of utterances and systems in each test set.
@@ 4.2 Exploration on Train Set Utilization
In audio quality assessment, whether for speech or singing, it is challenging to collect a sufficiently large and diverse dataset under a unified annotation standard. This makes it crucial to effectively leverage data annotated with different criteria. Following the approaches in AlignNet [10] and SHEET [8], we investigate two strategies on a self-supervised learning backbone. The first strategy is multi-dataset finetuning (MDF), in which we first train on the training set of the first annotation batch for 10 epochs and then continue finetuning on the entire training set. The second strategy is the use of the domain id, which corresponds to the specific annotation batch index.
As shown in Table 2, the SSL model without any auxiliary design and the SSL model equipped with both MDF and domain id achieve the best weighted SRCC at the system level and utterance level, respectively. Adding domain id leads to clear improvements on test3, suggesting that domain identifiers are particularly beneficial for test sets with fewer samples, as they help the model learn to mitigate domain discrepancies. Also, introducing MDF consistently improves performance on most metrics of test2, indicating that multi-dataset finetuning alleviates the confusion caused by different annotation standards across batches. Considering the weighted results across all test sets, the joint use of domain ids and MDF yields the most effective overall performance, while directly using the plain SSL model also proves to be a reasonable and convenient choice.
4.3 Exploration with Singing Quality Assessment
@@
Since SQA requires greater emphasis on melodic naturalness compared to speech assessment, we further explore several strategies for using pitch information. These strategies include the pitch histogram proposed in [26], as well as MIDI pitch and MIDI pitch variance, which are extracted from the raw waveform and quantized by $f_{0}$-to-MIDI conversion. As shown in the last three rows of Table 3, the pitch histogram yields slightly better performance than MIDI pitch, but the overall improvement over the SSL baseline remains marginal, underscoring the need for further exploration of how melodic cues can be more effectively integrated. Beyond these approaches, exploring how to incorporate melody scores and lyric scores into SQA constitutes a valuable avenue for future work.